After 1885 in a hospital, the number of births increase is shown with this equation- N=150(2^t/15) where n represents the number of births and t is the number of years since 1885. In what year would the annual birth tae in the hospital reach 3840?
Answer by Boreal(15235)

3840=150(2^(t/15)
divide both sides by 150
25.6=2^(t/15)
take log2 of both sides
log2 25.6=t/15
find log 2 (25.6) using change of base
log 10(25.6)/log 10(2)=4.678
That equals t/15
t=70.17 years or 1955 and it will exceed it in early 1956.